# Oblack Owhite m #3

### v1

@@ 5 Iterative Influence-Based Filtering for Online RL Training
Standard online RL algorithms typically treat all collected experiences uniformly. However, as our analysis in Sec. 4.1 has shown, some records can be harmful for learning. This likely contributes to the notorious sample inefficiency of online RL, a challenge widely acknowledged (Yu, 2018). Given this, a natural question arises: can we leverage the local data attribution framework to tackle this challenge?
We propose Iterative Influence-Based Filtering (IIF), building on the single-round interventions in Sec. 4.3. IIF filters records based on their computed influence scores, uses the resulting improved policy to sample new data, and repeats the cycle. This creates a loop for iterative refinement. We detail the algorithm below and showcase its effectiveness in traditional RL environments and RLHF for LLMs.
@@ 2 Function Update(model):
       $B\leftarrow$ $\mathcal{E}$, model, $n$)
@@
6      return PPOUpdate(model, ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{\text{filtered}}}$)
7for $\text{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}iter}=1$ to $T$ do
[…]
@@
Alg. 1 outlines IIF. Compared to standard PPO, IIF introduces an additional step of filtering (in red) between data collection and training. We further highlight the desiderata and IIF’s design choices.
@@
IIF employs a hyperparameter, $p$, which determines the amount of records to discard. We evaluate various $p$’s and report the best in Fig. 6. We observe that removing all negative-influence records ($p=100\%$) as in Wang et al. (2025a) is often suboptimal, likely due to the non-additivity of sample influence (Hu et al., 2024). Full ablation and recommendations for the choice of $p$ are in Sec. C.6.
@@ 5.2 Experiments in traditional RL environments
Experimental setup.   We evaluate IIF on the diverse set of RL environments introduced in Sec. 4.
Baselines: We compare IIF with standard PPO and a random filtering baseline (dropping a similar fraction of records). We additionally investigate an advantage based filtering heuristic in Sec. C.4 motivated by the characterization of bottom records in Sec. 4.1, as well as a TD error based heuristic in Sec. C.5 inspired by the Prioritized Experience Replay algorithm (Schaul et al., 2016).
Metrics: We quantify sample efficiency by the reduction in training rounds required for IIF to match standard training. For a performance level $v$ (measured by test return), let $m_{\text{std}}(v)$ and $m_{\text{IIF}}(v)$ be the earliest training rounds where standard training and IIF achieve performance at least $v$, respectively. The reduction at $v$ is defined as $(1-\nicefrac{{m_{\text{IIF}}(v)}}{{m_{\text{std}}(v)}})\times 100\%$. We report two metrics: $SE_{\text{ave}}$, the mean reduction over a list of strictly increasing performance levels reached by standard training, and $SE_{\text{peak}}$, the reduction at its peak. We measure computational cost by runtime; we similarly define $RT_{\text{peak}}$ as the reduction of runtime at the performance peak. Model performance is measured by the average test return over multiple episodes. See Sec. B.2 for further details on experimental setups.
@@
$SE_{\text{ave}}\ (\uparrow)$ 33.7% $\pm$ 3.4% 36.7% $\pm$ 6.5% 67.5% $\pm$ 5.1% 37.7% $\pm$ 6.1% 28.2% $\pm$ 1.3% 31.0% $\pm$ 8.7%
$SE_{\text{peak}}\ (\uparrow)$ 19.6% $\pm$ 8.4% 48.5% $\pm$ 0.8% 67.4% $\pm$ 4.4% 55.1% $\pm$ 2.9% 41.6% $\pm$ 5.7% 26.2% $\pm$ 8.0%
$RT_{\text{peak}}\ (\uparrow)$ 31.3% $\pm$ 3.3% 55.2% $\pm$ 1.0% 71.6% $\pm$ 1.1% 59.9% $\pm$ 0.7% 44.1% $\pm$ 2.3% 29.2% $\pm$ 0.7%
Figure 6: (a) Test returns over rounds for IIF vs. baselines. IIF speeds up learning and improves performance. Results are averaged over 3 random seeds. For Acrobot, we omit early rounds where returns rise from -500 to -200 for better visualization. (b) Sample efficiency and runtime metrics.
Results.   Fig. 6(a) presents the test returns for each environment; Fig. 6(b) summarizes the efficiency and runtime metrics. We report a detailed breakdown of runtime in Sec. C.7. Our key findings are summarized as follows: 1) IIF achieves substantial sample efficiency gains, showing a 20-67% reduction in training rounds required to match the standard training performance across environments. 2) The computational overhead of IIF is negligible, and offset by the reduced optimization time (see Sec. C.7), leading to significant improvement in runtime. 3) IIF’s final performance exceeds standard training in almost every environment. These observed gains stem from effective data attribution rather than mere data reduction—random filtering performs significantly worse than original training.
@@ 5.3 Extending IIF to RLHF for large language models
As the final part, we apply IIF to improve Reinforcement Learning from Human Feedback (RLHF). [^1] Another line of work focuses on improving reward modeling in RLHF (the stage before PPO) via preference data selection (Muldrew et al., 2024; Das et al., 2024; Shen et al., 2025); this is orthogonal to our work. Compared to standard PPO, RLHF introduces several key differences. First, the atomic unit shifts from state-action records to prompt-generation pairs, where each generation is a trajectory (or sequence) of tokens. Second, RLHF incorporates dual reward sources: a reward model evaluating the final generation, and a per-token KL divergence penalty to constrain deviation from a reference model.
@@
$f^{\text{seq}}(\theta)=\mathbb{E}_{(x,y)\sim\pi^{\text{ref}}}\left[\log\pi_{\theta}(y\mid x)\hat{A}^{\text{ref}}_{-1}(x,y)\right],$
where $x$ is the prompt, $y$ the generation, $\log\pi_{\theta}(y|x)=\sum_{i}\log\pi_{\theta}(y_{i}|x,y_{0},\ldots,y_{i-1})$ the log-probability of the sequence $y$ given $x$, and $\hat{A}^{\text{ref}}_{-1}$ the advantage estimate at the last token. This objective places emphasis on the reward model’s feedback at the last token of the sequence.
(a) Training Reward ($\uparrow$)
Figure 7: IIF improves the efficiency and performance of RLHF.
Experimental results: toxicity mitigation.   We consider the task of detoxifying LLMs using RLHF (Hugging Face, 2023), using gpt-neo-2.7B (Black et al., 2021) as our base model. Fig. 7 illustrates the effectiveness of our approach. We defer detailed experimental setups to Sec. B.3 and additional results (e.g., comparisons with using the target function $f^{\text{return}}$) in Sec. C.8.
We further highlight IIF’s substantial gains in computational efficiency. In RLHF, sample efficiency is typically not a major concern since interacting with the environment is essentially inference, which is lightweight comparatively. The dominant cost lies in training. IIF filters out negative-influence records ($\sim$50of data), effectively halving the optimization time per round. Furthermore, IIF accelerates learning, requiring less than half the number of rounds to surpass standard training. The overhead of influence calculation is minimal. Collectively, these factors result in an approximate 4$\times$ reduction in total runtime. A detailed runtime breakdown is provided in Sec. C.9.
@@ 6 Conclusion and Limitations
Counterfactual interpretation.   Finally, our local attribution framework, while powerful, lacks a clear counterfactual interpretation. This limitation partly stems from TracIn itself, but primarily from the fundamental difficulty of tracking causal effects across the circular data-model dependency inherent in online RL, as discussed in Sec. 3. We encourage future work to tackle this open problem.
@@ Acknowledgements
YH and HZ are partially supported by an NSF IIS grant with No. 2416897, and an NVIDIA Academic Grant Program. HZ would like to thank the support from a Google Research Scholar Award.
Appendix A Related Work
Interpretability in reinforcement learning has become a central research theme because real-world deployment requires agents that are trustworthy and reliable [Arulkumaran et al., 2017, Sutton and Barto, 2018, Milani et al., 2024, Cheng et al., 2025]. Early studies emphasize feature-level explanations: they highlight regions of the observation space that most influence an agent’s decisions, often through saliency maps or attention heatmaps [Zahavy et al., 2016, Greydanus et al., 2018, Iyer et al., 2018, Mott et al., 2019, Atrey et al., 2020, Puri et al., 2020]. A complementary thread seeks policy-level explanations. These works approximate learned policies with human-interpretable rules [Verma et al., 2018, Soares et al., 2020], design transparent architectures [Topin et al., 2021, Demircan et al., 2025], or dissect reward functions to clarify action choices [Juozapaitis et al., 2019, Liu and Zhu, 2025]. More recently, researchers have probed how entire training trajectories shape behavior [Deshmukh et al., 2023].
Zooming in further, identifying critical states offers a finer-grained view of decision making. Several approaches address offline settings [Guo et al., 2021, Yu et al., 2023, Liu et al., 2023, Rishav et al., 2025]. Closer to our focus are methods that target online RL such as lazy-MDP [Jacq et al., 2022], StateMask [Cheng et al., 2023] and RICE [Cheng et al., 2024]. Lazy-MDP augments the action space with a “lazy” action and penalizes non-lazy choices; states where the agent still acts are interpreted as important. However, this approach requires modifying the training pipeline. StateMask and RICE train an auxiliary mask network alongside the policy, forcing random actions in selected states while keeping returns roughly unchanged; masked states are deemed non-critical. Nevertheless, these methods crucially rely on the policy being sufficiently developed, which limits their applicability when agents are still learning in complex environments.
Moving beyond these constraints, our work introduces data attribution as a principled lens for interpretability in online RL. This approach not only closes a key methodological gap in the literature but also delivers fresh insights for RL researchers and practitioners, and informs more efficient and effective training.
Appendix B Detailed Experimental Setups
B.1 Standard RL Environments
We offer a detailed description of the RL environments used in our experiments in Table 1.
@@ B.1 Standard RL Environments
FrozenLake FrozenLake-v1333https://gymnasium.farama.org/environments/toy_text/frozen_lake/, map=4x4,  slippery=False Navigate from start to goal without falling into holes 1 discrete integer: agent position index on the grid 4 discrete actions: {Left, Down, Right, Up} Sparse: +1 on reaching goal, 0 otherwise
Acrobot Acrobot-v1444https://gymnasium.farama.org/environments/classic_control/acrobot/ Swing up the link to reach a target height $\mathbb{R}^{6}$, providing information about the two rotational joint angles and their angular velocities 3 discrete actions: $\{-1,0,1\}$ torque (N m) Dense: -1 per step until reaching the target height
Highway    [Leurent, 2018] highway-v0555https://highway-env.farama.org/environments/highway/, vehicle_count= 10 Drive at high speed while avoiding collisions Kinematic Observation: $5\times 5$ array of ego and nearby vehicles, including their location and speed 5 discrete actions: {LANE_LEFT, IDLE, LANE_RIGHT, FASTER, SLOWER} Dense: $\nicefrac{{(v-v_{\min})}}{{(v_{\max}-v_{\min})}}$$-b\cdot\mathrm{collision}$ at each step
LunarLander LunarLander-v2666https://gymnasium.farama.org/environments/box2d/lunar_lander/ Land safely on the pad from flight $\mathbb{R}^{8}$: the coordinates of the lander, its linear velocities, angle, angular velocity, and whether each leg is in contact with the ground 4 discrete actions: {do nothing, fire left, fire main, fire right} Dense: +10 per leg contact; –0.03 per side-engine step; –0.3 per main-engine step; +100 on safe landing; –100 on crash; distance/velocity/angle terms
BipedalWalker BipedalWalker-v3777https://gymnasium.farama.org/environments/box2d/bipedal_walker/ Traverse rough terrain without falling $\mathbb{R}^{24}$: hull angle speed, angular velocity, horizontal & vertical speed, joints positions & angular speed, legs contact with ground, 10 lidar measurements 4 continuous actions: motor speed values in $[-1,1]$ for 4 joints at hips and knees Dense: +1 per forward step; -100 on fall; small penalty proportional to torque magnitude
B.2 Experimental setups for standard RL
@@
We adopt Stable-Baselines3888https://stable-baselines3.readthedocs.io/en/master/index.html [Raffin et al., 2021] (MIT license) as our training framework for the standard RL experiments. We use PPO [Schulman et al., 2017] as our RL algorithm and adopt the default training hyperparamters and network architectures for most environments unless otherwise specified.
@@
B.3 Experimental setups for RLHF
@@
We adopt TRL [^9] https://huggingface.co/docs/trl/index [von Werra et al., 2020] (Apache-2.0 license) as our training framework to fine-tune the based model via PPO. The dataset for PPO training is real-toxicity-prompts [^10] https://huggingface.co/datasets/allenai/real-toxicity-prompts [Gehman et al., 2020] (Apache-2.0 license). For each example, we extract the first 10-15 tokens as a prompt, generate a 30-token continuation, and score it with the reward model, a toxicity detector LFTW R4 Target [^11] https://huggingface.co/facebook/roberta-hate-speech-dynabench-r4-target[Vidgen et al., 2021]. The reward signal is the raw logits of the label “neutral” of the detector.
@@
Appendix C Additional Experimental Results
C.1 More demonstrations of harmful records

### v2

@@ 5 Iterative Influence-Based Filtering for Online RL Training
Standard online RL algorithms typically treat all collected experiences uniformly. However, as our analysis in Sec.˜4.1 has shown, some records can be harmful for learning. This likely contributes to the notorious sample inefficiency of online RL, a challenge widely acknowledged (Yu, 2018). Given this, a natural question arises: can we leverage the local data attribution framework to tackle this challenge?
We propose Iterative Influence-Based Filtering (IIF), building on the single-round interventions in Sec.˜4.3. IIF filters records based on their computed influence scores, uses the resulting improved policy to sample new data, and repeats the cycle. This creates a loop for iterative refinement. We detail the algorithm below and showcase its effectiveness in traditional RL environments and RLHF for LLMs.
@@ 2 Function Update(model):
  $B\leftarrow$ CollectTransitions($\mathcal{E}$, model, $n$)
@@
6 return PPOUpdate(model, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B_{\emph{filtered}}$})
7for $\emph{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}iter}=1$ to $T$ do
8  $\textnormal{{model}}\;\leftarrow\;\textnormal{{Update}}(\textnormal{{model}})$
@@
Alg.˜1 outlines IIF. Compared to standard PPO, IIF introduces an additional step of filtering (in red) between data collection and training. We further highlight the desiderata and IIF’s design choices.
@@
IIF employs a hyperparameter, $p$, which determines the amount of records to discard. We evaluate various $p$’s and report the best in Fig.˜6. We observe that removing all negative-influence records ($p=100\%$) as in Wang et al. (2025a) is often suboptimal, likely due to the non-additivity of sample influence (Hu et al., 2024). Full ablation and recommendations for the choice of $p$ are in Sec.˜B.6.
@@ 5.2 Experiments in traditional RL environments
Experimental setup.   We evaluate IIF on the diverse set of RL environments introduced in Sec.˜4.
Baselines: We compare IIF with standard PPO and a random filtering baseline (dropping a similar fraction of records). We additionally investigate an advantage based filtering heuristic in Sec.˜B.4 motivated by the characterization of bottom records in Sec.˜4.1, as well as a TD error based heuristic in Sec.˜B.5 inspired by the Prioritized Experience Replay algorithm (Schaul et al., 2016).
Metrics: We quantify sample efficiency by the reduction in training rounds required for IIF to match standard training. For a performance level $v$ (measured by test return), let $m_{\text{std}}(v)$ and $m_{\text{IIF}}(v)$ be the earliest training rounds where standard training and IIF achieve performance at least $v$, respectively. The reduction at $v$ is defined as $(1-\nicefrac{{m_{\text{IIF}}(v)}}{{m_{\text{std}}(v)}})\times 100\%$. We report two metrics: $SE_{\text{ave}}$, the mean reduction over a list of strictly increasing performance levels reached by standard training, and $SE_{\text{peak}}$, the reduction at its peak. We measure computational cost by runtime; we similarly define $RT_{\text{peak}}$ as the reduction of runtime at the performance peak. Model performance is measured by the average test return over multiple episodes. See Sec.˜A.2 for further details on experimental setups.
@@
$SE_{\text{ave}}\ (\uparrow)$
34.0% $\pm$ 2.0% 36.7% $\pm$ 6.5% 65.8% $\pm$ 3.3% 37.7% $\pm$ 6.1% 26.0% $\pm$ 1.8% 31.0% $\pm$ 8.7%
$SE_{\text{peak}}\ (\uparrow)$
19.2% $\pm$ 5.9% 48.5% $\pm$ 0.8% 61.7% $\pm$ 4.1% 55.1% $\pm$ 2.9% 39.7% $\pm$ 3.7% 26.2% $\pm$ 8.0%
$RT_{\text{peak}}\ (\uparrow)$ 29.5% $\pm$ 2.9% 55.2% $\pm$ 1.0% 69.1% $\pm$ 1.7% 59.9% $\pm$ 0.7% 44.9% $\pm$ 2.5% 29.2% $\pm$ 0.7%
Figure 6: (a) Test returns over rounds for IIF vs. baselines. IIF speeds up learning and improves performance. Results are averaged over 5 random seeds. For Acrobot, we omit early rounds where returns rise from -500 to -200 for better visualization. (b) Sample efficiency and runtime metrics.
Results.   Fig.˜6(a) presents the test returns for each environment; Fig.˜6(b) summarizes the efficiency and runtime metrics. We report a detailed breakdown of runtime in Sec.˜B.9. Our key findings are summarized as follows: 1) IIF achieves substantial sample efficiency gains, showing a 20-67% reduction in training rounds required to match the standard training performance across environments. 2) The computational overhead of IIF is negligible, and offset by the reduced optimization time (see Sec.˜B.9), leading to significant improvement in runtime. 3) IIF’s final performance exceeds standard training in almost every environment. These observed gains stem from effective data attribution rather than mere data reduction: random filtering performs significantly worse than original training.
@@ 5.3 Extending IIF to RLHF for large language models
As the final part, we apply IIF to improve Reinforcement Learning from Human Feedback (RLHF). [^2] Another line of work focuses on improving reward modeling in RLHF (the stage before PPO) via preference data selection (Muldrew et al., 2024; Das et al., 2024; Shen et al., 2025); this is orthogonal to our work. Compared to standard PPO, RLHF introduces several key differences. First, the atomic unit shifts
(a) Training Reward ($\uparrow$)
Round
Round
Figure 7: IIF improves the efficiency and performance of RLHF.
from state-action records to prompt-generation pairs, where each generation is a trajectory (or sequence) of tokens. Second, RLHF incorporates dual reward sources: a reward model evaluating the final generation, and a per-token KL divergence penalty to constrain deviation from a reference model.
@@
$f^{\text{seq}}(\theta)=\mathbb{E}_{x\sim D_{\text{val}},y\sim\pi^{\text{ref}}(\cdot\mid x)}\left[\log\pi_{\theta}(y\mid x)\hat{A}^{\text{ref}}_{-1}(x,y)\right],$
where $x$ is a prompt drawn from the validation set $D_{\text{val}}$, $y$ the generation, $\log\pi_{\theta}(y|x)=\sum_{i}\log\pi_{\theta}(y_{i}|x,y_{0},\ldots,y_{i-1})$ the log-probability of the sequence $y$ given $x$, and $\hat{A}^{\text{ref}}_{-1}$ the advantage estimate at the last token. This objective emphasizes the reward model’s feedback at the last token.
Experimental results: toxicity mitigation.   We consider the task of detoxifying LLMs using RLHF (Hugging Face, 2023), using gpt-neo-2.7B (Black et al., 2021) as our base model. Fig.˜7 illustrates the effectiveness of our approach. We defer detailed experimental setups to Sec.˜A.3 and additional results (e.g., comparisons with using the target function $f^{\text{return}}$) in Sec.˜B.11.
[…]
6 Related Work
Interpretability in reinforcement learning has become a central research theme because real‑world deployment requires agents that are trustworthy and reliable (Arulkumaran et al., 2017; Sutton and Barto, 2018; Milani et al., 2024; Cheng et al., 2025). Early studies emphasize feature-level explanations: they highlight regions of the observation space that most influence an agent’s decisions, often through saliency maps or attention heatmaps (Zahavy et al., 2016; Greydanus et al., 2018; Mott et al., 2019; Atrey et al., 2020; Puri et al., 2020). A complementary thread seeks policy-level explanations. These works approximate learned policies with human-interpretable rules (Verma et al., 2018; Soares et al., 2020), design transparent architectures (Topin et al., 2021; Demircan et al., 2025), or dissect reward functions to clarify action choices (Juozapaitis et al., 2019; Liu and Zhu, 2025). More recently, researchers have probed how entire training trajectories shape behavior (Deshmukh et al., 2023).
Zooming in further, identifying critical states offers a finer‑grained view of decision making. Several approaches address offline settings (Guo et al., 2021; Yu et al., 2023; Liu et al., 2023; Rishav et al., 2025). Closer to our focus are methods that target online RL such as lazy-MDP (Jacq et al., 2022), StateMask (Cheng et al., 2023) and RICE (Cheng et al., 2024). Lazy-MDP augments the action space with a “lazy” action and penalizes non‑lazy choices; states where the agent still acts are interpreted as important. However, this approach requires modifying the training pipeline. StateMask and RICE train an auxiliary mask network alongside the policy, forcing random actions in selected states while keeping returns roughly unchanged; masked states are deemed non‑critical. Nevertheless, these methods crucially rely on the policy being sufficiently developed, which limits their applicability when agents are still learning in complex environments.
Moving beyond these constraints, our work introduces data attribution as a principled lens for interpretability in online RL. This approach closes a key methodological gap in the literature, delivers fresh insights for RL researchers and practitioners, and informs more efficient and effective training.
@@ 7 Conclusion and Limitations
Counterfactual interpretation.   Finally, our local attribution framework, while powerful, lacks a clear counterfactual interpretation. This limitation partly stems from TracIn itself, but primarily from the fundamental difficulty of tracking causal effects across the circular data-model dependency inherent in online RL, as discussed in Sec.˜3. We encourage future work to tackle this open problem.
@@ Acknowledgements
We thank the anonymous NeurIPS 2025 reviewers for their constructive feedback. YH thanks Haozhe Si for assistance in setting up an NVIDIA instance. YH and HZ are partially supported by NSF IIS Grant No.2416897 and the NVIDIA Academic Grant Program. HZ also acknowledges support from a Google Research Scholar Award.
Appendix A Detailed Experimental Setups
A.1 Standard RL environments
We offer a detailed description of the RL environments used in our experiments in Table˜1.
@@ A.1 Standard RL environments
FrozenLake FrozenLake-v1444https://gymnasium.farama.org/environments/toy_text/frozen_lake/, map=4x4,  slippery=False Navigate from start to goal without falling into holes 1 discrete integer: agent position index on the grid 4 discrete actions: {Left, Down, Right, Up} Sparse: +1 on reaching goal, 0 otherwise
Acrobot Acrobot-v1555https://gymnasium.farama.org/environments/classic_control/acrobot/ Swing up the link to reach a target height $\mathbb{R}^{6}$, providing information about the two rotational joint angles and their angular velocities 3 discrete actions: $\{-1,0,1\}$ torque (N m) Dense: -1 per step until reaching the target height
Highway    [Leurent, 2018] highway-v0666https://highway-env.farama.org/environments/highway/, vehicle_count= 10 Drive at high speed while avoiding collisions Kinematic Observation: $5\times 5$ array of ego and nearby vehicles, including their location and speed 5 discrete actions: {LANE_LEFT, IDLE, LANE_RIGHT, FASTER, SLOWER} Dense: $\nicefrac{{(v-v_{\min})}}{{(v_{\max}-v_{\min})}}$$b\cdot\mathrm{collision}$ at each step
LunarLander LunarLander-v2777https://gymnasium.farama.org/environments/box2d/lunar_lander/ Land safely on the pad from flight $\mathbb{R}^{8}$: the coordinates of the lander, its linear velocities, angle, angular velocity, and whether each leg is in contact with the ground 4 discrete actions: {do nothing, fire left, fire main, fire right} Dense: +10 per leg contact; –0.03 per side‑engine step; –0.3 per main‑engine step; +100 on safe landing; –100 on crash; distance/velocity/angle terms
BipedalWalker BipedalWalker-v3888https://gymnasium.farama.org/environments/box2d/bipedal_walker/ Traverse rough terrain without falling $\mathbb{R}^{24}$: hull angle speed, angular velocity, horizontal & vertical speed, joints positions & angular speed, legs contact with ground, 10 lidar measurements 4 continuous actions: motor speed values in [1,1][-1,1] for 4 joints at hips and knees Dense: +1 per forward step; -100 on fall; small penalty proportional to torque magnitude
A.2 Experimental setups for standard RL
@@
We adopt Stable-Baselines3999https://stable-baselines3.readthedocs.io/en/master/index.html [Raffin et al., 2021] (MIT license) as our training framework for the standard RL experiments. We use PPO [Schulman et al., 2017] as our RL algorithm and adopt the default training hyperparamters and network architectures for most environments unless otherwise specified.
@@
A.3 Experimental setups for RLHF
@@
We adopt TRL [^10] https://huggingface.co/docs/trl/index [von Werra et al., 2020] (Apache-2.0 license) as our training framework to fine-tune the based model via PPO. We employ LoRA [Hu et al., 2022] to perform PEFT fine-tuning, with a rank of 16, $\alpha$ of 32 and dropout of 0.05. The dataset for PPO training is real-toxicity-prompts [^11] https://huggingface.co/datasets/allenai/real-toxicity-prompts [Gehman et al., 2020] (Apache-2.0 license). For each example, we extract the first 10-15 tokens as a prompt, generate a 30-token continuation, and score it with the reward model, a toxicity detector LFTW R4 Target [^12] https://huggingface.co/facebook/roberta-hate-speech-dynabench-r4-target[Vidgen et al., 2021]. The reward signal is the raw logits of the label “neutral” of the detector.
@@
Appendix B Additional Experimental Results
B.1 More demonstrations of harmful records